# Longitudinal Abuse and Sentiment Analysis of Hollywood Movie Dialogues using LLMs

### v2

@@ Abstract
Over the past decades, there has been an increasing concern about the prevalence of abusive and violent content in Hollywood movies. This study uses Large Language Models (LLMs) to explore the longitudinal abuse and sentiment analysis of Hollywood Oscar and blockbuster movie dialogues from 1950 to 2024. By employing fine-tuned LLMs, we analyse subtitles for over a thousand movies categorised into four genres to examine the trends and shifts in emotional and abusive content over the past seven decades. Our findings reveal significant temporal changes in movie dialogues, which reflect broader social and cultural influences. Overall, the emotional tendencies in the films are diverse, and the detection of abusive content also exhibits significant fluctuations. The results show a gradual rise in abusive content in recent decades, reflecting social norms and regulatory policy changes. Genres such as thrillers still present a higher frequency of abusive content that emphasises the ongoing narrative role of violence and conflict. At the same time, underlying positive emotions such as humour and optimism remain prevalent in most of the movies. Furthermore, the gradual increase of abusive content in movie dialogues has been significant over the last two decades, where Oscar-nominated movies overtook the top ten blockbusters.
@@
Large Language Models, Sentiment Analysis, Abusive Detection, Hollywood, Oscar, Blockbusters, Dialogues
journal: .
[…]
[inst1]organization=Transitional Artificial Intelligence Research Group, School of Mathematics and Statistics, University of New South Wales, city=Kensington, state=Sydney, country=Australia
[…]
[inst2]organization=Centre for Artificial Intelligence and Innovation, Pingla Institute, state=Sydney, country=Australia
@@ 1 Introduction
The cinema has been present in various forms for over a century, serving multiple purposes as an artistic medium through movies and series [1]. Since the age of television [2], the cinema has evolved from a mere source of entertainment to an influential tool that includes politics [3], fashion, and cultural and social changes [4], supporting activism and policy changes, e.g., advocacy for queer activism and same-sex marriage [5]. Furthermore, there has also been activism with movies about climate change [6] including natural disasters (e.g. Day After Tomorrow [7]) and dystopian futures showing a water crisis (e.g. Mad Max [8]).
@@
The cinema had profound positive and negative cultural and social effects. Some of the negative impacts are an increase in crime rate [9], such as drug use and distribution [10], and sex crimes [11]. Violence is an important narrative tool for cinema, despite some of its adverse social effects. Violence and abuse in cinema have been used to enhance a viewer’s experience, and boost movie profits [12]; therefore, the prevalence of violence in movies has significantly escalated over the past few decades [13]. However, the relationship between movies and human behaviour is of utmost significance [14] and violent movies often have an inciting effect [15]. Violent movies promote aggression and lead to antisocial behaviour [16] that includes aggressive behaviour as well as cognitive, social, and emotional problems, often accompanied by a lack of empathy [17]. The cinema, including movies, series, and documentary media content, has lasting emotional effects on young viewers. Canter et al. [18] reported that different people have different emotional and behavioural reactions to violent movies. Moreover, they reported that children’s exposure to cinema and related media can lead to increased fear and anxiety.
The way of obtaining entertainment has become more diverse with the Internet and emerging technologies. OTT (Over-the-top) [19] is a digital media platform for film and television content delivered through the Internet. OTT service providers such as Netflix, Amazon Prime Video and Disney+ use a subscription system or pay-per-view to provide users with a huge amount of movie and television resources, which has become an important channel for current entertainment consumption [20]. The future could have OTT platforms gaining the upper hand over traditional movie theatres due to factors, such as diverse content, language availability, ease of viewing, and affordable costs [21]. As OTT platforms become one of the main channels for entertainment, we can also examine the challenges that may arise.
The field of Natural Language Processing (NLP) [22] lies at the confluence of artificial intelligence and linguistics, involving the design and implementation of algorithms that interact with human language using methods such as deep learning [23]. Over the past few years, the widespread use of deep learning models has propelled NLP, enabling it to deal efficiently with more complex language problems [23]. Large language models (LLMs) [24] are deep learning models with an enormous number of parameters trained in an unsupervised manner using a large corpus of text data [24]. LLMs have demonstrated impressive abilities in NLP tasks [25] such as sentiment analysis, question answering, text summarising, and language translation. Sentiment analysis [26] using deep learning models [27], focuses on detecting human emotions in text such as ’anxious’, ’happy’, ’optimistic’ and sentiment polarity prediction such as positive, negative, or neutral opinions. Sentiment analysis has extensive applications in different domains, encompassing social media analytics, public sentiment monitoring, product review evaluation, and customer feedback assessment [28, 26].
Recurrent neural networks (RNNs) [29] are deep learning models that possess the ability to acquire features and capture long-term dependencies from sequential and time-series data [29]. The Transformer model is a deep learning model that features an encoder-decoder architecture and attention-based mechanism inspired by cognitive science [30]. Although RNNs have been widely used, the introduction of the Transformer model has marked a significant shift in NLP, paving the way for pre-trained large language models. Recent innovations in pre-trained language models, also known as LLMs [31], include BERT (Bidirectional Encoder Representation from Transformers) [32] and GPT (Generative Pre-trained Transformers) [31], which have been prominent in a wide range of NLP tasks with implementations including ChatGPT, GPT-4o, Gemini, and Grok [33].
@@
Abuse detection using NLP and LLMs has been prominent due to the needs of social media, where real-time management of discussions are needed to avoid and limit cyber-bullying. Caselli et al. [34] introduced the Hate-BERT model for abusive content detection, achieving better accuracy in detecting various abusive content categories including offensive content, abusive content, and hateful language. Park et al. [35] introduced three methods: debiased word embeddings, gender swap data augmentation, and fine-tuning with a larger corpus to reduce gender biases in abuse detection. Nobata et al. [36] effectively developed a machine learning-based approach to detect hate speech in online user comments. Mozafari et al. [37] added a Bidirectional-LSTM layer to BERT’s architecture to enhance its classification capabilities. Mnassri et al. [38] explored various integrated approaches to improve BERT for hate speech detection, and reported that combining BERT with other deep learning models improved accuracy.
Furthermore, BERT and GPT use the Transformer model [39] to leverage for different tasks in NLP. GPT uses the previous context information in a text sequence to predict words (i.e., unidirectional), making the model simpler, computationally faster, and more suitable for natural language generation tasks [40]. BERT is specifically designed to pre-train deep bidirectional representations from unlabelled text, effectively incorporating both the left and right context of a sentence. The pre-trained BERT model can be fine-tuned with a single additional output layer to achieve state-of-the-art performance in various tasks [32]. Therefore, BERT is particularly suitable for this study, as it allows effective sentiment analysis of movie dialogues while detecting violent and abusive content.
@@
The defining characteristic of longitudinal analysis lies in the repeated measurements taken on the same individuals or subject o study over time, enabling a direct examination of temporal changes [41]. The primary objective of a longitudinal study is to delineate the dynamics of response variation and identify influential factors [41]. By capturing within-individual changes through multiple assessments, one can effectively examine temporal variations [41].
@@
In this study, we present a framework for abuse detection and sentiment analysis on dialogues from selected movies. We select Hollywood movies (films) spanning from 1950 to 2024 based on Oscar nominations and box-office collections. We analyze decadal changes in the levels of abuse and violence, utilising sentiment analysis across major categories, including Oscar nominations and blockbuster hits. We further classify the selected movies into four categories to provide an analysis of how the abuse and sentiment analysis impact different categories of movies, e.g. action vs drama. We use an abuse detection dataset [34] to fine-tune LLMs (BERT-based model) and provide a longitudinal study. Our framework also provides sentiment analysis, where we use the SenWave [42] dataset to fine-tune the LLM. We also curate the subtitle dataset of more than a thousand movies for further analysis.
@@
The structure of our research is as follows: Section 2 provides background information, and Section 3 discusses the methodology, including the framework, data processing, and modelling. In Section 4, we present and analyze the results. Section 5 offers a discussion and outlines directions for further research. Finally, Section 6 concludes the study.
@@ 2.1 Cinema: impact on society
Film as a medium is not only considered one of the most important sources of entertainment, but also a powerful tool for influencing society and culture [43]. Although the themes of love, friendship, and family relationships in films inspire audiences positively, some films with explicit, vulgar, or violent content can negatively impact viewers’ psychological well-being [44], affecting children and adolescents [45] . Furthermore, the characteristics of the protagonist, such as justice, wisdom and bravery, have a positive influence on the growth of young people [46].
The United Nations Educational, Scientific and Cultural Organisation (UNESCO) [47] presented a comprehensive study of the impact of media violence on the audience. Most studies show that the relationship between media abuse and “real” abuse is interactive, and the perpetuation and persistence of media violence fuelling a global culture of aggression [48, 49, 50]. The American psychologist, Albert Bandura [51] reported that when people consciously or unconsciously remember and preserve abusive content, and unconsciously activate them when there are setbacks in social interactions. According to the American Academy of Pediatrics [52], television violence accounts for 10 percent of the sources of learning about violent behaviour in children. Vitasari [53] reported that the use of abusive words is now more often used in films. The uninhibited display of emotions is an indication of cultural transformation, resulting in a rise in the utilisation of profane language. Hasibuan [54] contended that many films now often use swearing in dialogues to make them more interesting. Novari [55] reported that characters in movies do not always use swearing to curse but to express their other emotions and feelings in different situations. Baid et al. [56] demonstrated that morphing learning from the social media domain can effectively classify hateful and offensive speech in movie subtitles. The authors used domain-adaptive and fine-tuned transfer learning techniques to an existing social media dataset. Mathew et al. [57] presented the HateXplain model based on BERT to classify movie subtitles, which achieved better accuracy than previous machine learning models [56]. Deep learning models have been prominent in sentiment analysis [58]; hence, we will use them to analyze movie dialogues.
@@ 2.2 Web series and movies
Web series [59] is a short-form interstitial content distributed through online platforms. Since the 1990s onwards, with the advent of technological change and the emergence of the new media market, web series have become prominent, splitting the market of limited television and film [60]. During COVID-19, as countries went into embargo and for reasons such as economic depression, it gave online sketches a huge opportunity to grow and expand more widely into the population. Traditional movies have been facing a huge challenge due to OTT platforms, and a study shows that as the length of movies increases, people expect movies released in theatres to be released simultaneously on OTT as well [61]. Therefore, web series have become one of the most popular trends on OTT platforms [62]. However, the content preferences for films and web series are very different. In a related study, Wadhwa et al. [63] conducted sentiment analysis for ten films of different genres as well as ten web series during 2017-2019. The authors accessed Twitter data related to each show was accessed through hashtags and 3000 tweets were counted to determine the sentiment of each web series and categorised into positive and negative sentiment. The results found that horror web series were more popular with viewers than horror films, but comedy, romance and crime films were more popular than web series.
@@ 3.1 Data
The Oscars (Academy Awards)  [^1] Oscar website: https://www.oscars.org/, presented annually by the Academy of Motion Picture Arts and Sciences (AMPAS) recognise artistic and technical achievements in movies predominately from the USA (Hollywood) [64]. Historically, the Oscars have nominated various films, including historical dramas, biopics, romances, science fiction, etc. Although the drama category plays a big part in the Academy Awards; horror, comedy and independent films have gradually gained more recognition in recent years with movies such as ”Get Out” and ”Backstage.” This diversity allows Oscar-nominated films to represent different periods and audiences’ aesthetic tastes [64]. The top 10 blockbuster films at the global box office each year represent the most popular movie genres and themes of the year. These films are often able to attract a large audience and have a wide impact on a global scale. The top 10 blockbuster films consist of genres such as action, science fiction, comedy, and animation, and offer insights into cultural preferences and emotional resonance. The popularity of these films in different cultural contexts reflects the general acceptance of their emotional expressions and scenes of violence. Therefore, we chose to study the subtitles of Oscar-nominated movies and the top 10 movie subtitles at the box office each year from 1950 to 2024.
@@
We obtained the top 10 Oscar-nominated and top 10 movies at the global Box Office for 1950 to 2024 from a website known as boxofficemojo  [^2] https://www.boxofficemojo.com/. We used Opensubtitles [^3] https://www.opensubtitles.com/en to extract the subtitles (dialogues) for the top 10 Oscar-nominated and global Box Office films from 1950 to 2024. Because the number of Oscar-nominated films varies from year to year, we obtained a total of 1026 film subtitles and grouped them by year and genre to analyze changes in violence and abuse content and emotional changes in films through time changes. The dataset is available via Kaggle  [^4] https://www.kaggle.com/datasets/mlopssss/subtitles and also documented in our GitHub repository  [^5] https://github.com/sydney-machine-learning/sentimentanalysis-Hollywood for this study.
@@
We categorise the movies into four broad categories, grouping similar categories into one according to the categories defined by the IMBD (International Movie Database)  [^6] https://www.imdb.com/ website to facilitate comparison of different movie categories. We group the categories of action, adventure, and crime as action, and group horror and thriller categories as thrillers. We keep the comedy and drama categories unchanged. Since most movies are defined by three category labels on the IMBD website, we take the first label as the main category of the movie as the basis for classification and report the number of movies under each category in Table 1.
@@
Table 1: Numbers of movies in the four categories that include drama, action, comedy and thriller.

### v3

# Longitudinal Abuse and Sentiment Analysis of Hollywood Movie Dialogues using Language Models

@@ Abstract
Over the past decades, there has been an increase in the prevalence of abusive and violent content in Hollywood movies. In this study, we use language models to explore the longitudinal abuse and sentiment analysis of Hollywood Oscar and blockbuster movie dialogues from 1950 to 2024. We provide an analysis of subtitles for over a thousand movies, which are categorised into four genres. We employ fine-tuned language models to examine the trends and shifts in emotional and abusive content over the past seven decades. Findings reveal significant temporal changes in movie dialogues, which reflect broader social and cultural influences. Overall, the emotional tendencies in the films are diverse, and the detection of abusive content also exhibits significant fluctuations. The results show a gradual rise in abusive content in recent decades, reflecting social norms and regulatory policy changes. Genres such as thrillers still present a higher frequency of abusive content that emphasises the ongoing narrative role of violence and conflict. At the same time, underlying positive emotions such as humour and optimism remain prevalent in most of the movies. Furthermore, the gradual increase of abusive content in movie dialogues has been significant over the last two decades, where Oscar-nominated movies overtook the top ten blockbusters.
@@
Language Models, Sentiment Analysis, Abusive Detection, Hollywood, Oscar, Blockbusters, Dialogues
@@ 1 Introduction
The cinema has been present in various forms for over a century, serving multiple purposes as an artistic medium through movies and series (Nowell-Smith,, 1996). Since the age of television (Buonanno and Radice,, 2008), the cinema has evolved from a mere source of entertainment to an influential tool that includes politics (Tzioumakis and Molloy,, 2016), fashion, and cultural and social changes (Kerrigan,, 2018), supporting activism and policy changes, e.g., advocacy for queer activism and same-sex marriage (Schoonover and Galt,, 2015). Furthermore, there has also been activism with movies about climate change (Sakellari,, 2015) including natural disasters (e.g. Day After Tomorrow (Leiserowitz,, 2004)) and dystopian futures showing a water crisis (e.g. Mad Max (Hassler-Forest,, 2017)). Although cinema has played an important role in shaping culture and inspiring activism, it has also been criticised for its potential negative influences, particularly through the depiction of violence and abusive content.
@@
The cinema had profound positive and negative cultural and social effects. Some of the negative impacts are an increase in crime rate (Young,, 2009), such as drug use and distribution (Grist,, 2013), and sex crimes (Gunasekera et al.,, 2005). Violence is a powerful narrative tool in cinema, despite its adverse social effects. Violence and abuse in cinema have been used to enhance a viewer’s experience, and boost movie profits (Barranco et al.,, 2020); therefore, the prevalence of violence in movies has significantly escalated over the past few decades (Markey et al.,, 2015). However, the relationship between movies and human behaviour is of utmost significance (Fearing,, 1947) and violent movies often have an inciting effect (Dunand et al.,, 1984). Violent movies promote aggression and lead to antisocial behaviour (Berkowitz et al.,, 1963) that includes aggressive behaviour as well as cognitive, social, and emotional problems, often accompanied by a lack of empathy (Paik and Comstock,, 1994). Anderson and Bushman (Anderson and Bushman,, 2002) presented a meta-analysis and reported strong links between exposure to violent media and increased aggression in viewers. Later, they provided a theoretical synthesis explaining how exposure to violent media affects aggression using three pathways: cognitive, affective, and arousal mechanisms (Anderson and Bushman,, 2018). The cinema includes movies, series, and documentaries, which can have lasting emotional effects on young viewers. Canter et al. (Cantor,, 2001) reported that different people have different emotional and behavioural reactions to violent movies. Moreover, they reported that children’s exposure to cinema and related media can lead to increased fear and anxiety. These social concerns are further amplified in the contemporary media landscape, where digital platforms have transformed how audiences consume cinematic content.
The way of obtaining entertainment has become more diverse with the Internet and emerging technologies. OTT (Over-the-top) (Varghese and Chinnaiah,, 2021) is a digital media platform for film and television content delivered through the Internet. OTT service providers such as Netflix, Amazon Prime Video and Disney+ use a subscription system or pay-per-view to provide users with a huge amount of movie and television resources, which has become an important channel for current entertainment consumption (Eklund,, 2023). The future could have OTT platforms gaining the upper hand over traditional movie theatres due to factors, such as diverse content, language availability, ease of viewing, and affordable costs (Gaonkar et al.,, 2022). As OTT platforms become one of the main channels for entertainment, The study will also examine the challenges that may arise. Given the scale and diversity of content available on these platforms, computational approaches such as Natural Language Processing (NLP) have become essential for systematically analysing large volumes of textual data, including film dialogues. The field of NLP (Cambria and White,, 2014) lies at the confluence of artificial intelligence and linguistics, involving the design and implementation of algorithms that interact with human language using methods such as deep learning (Otter et al.,, 2021). Over the past few years, the widespread use of deep learning models has propelled NLP, enabling it to deal efficiently with more complex language problems (Otter et al.,, 2021). Sentiment analysis (Medhat et al.,, 2014) using deep learning models (Ain et al.,, 2017), focuses on detecting human emotions in text, such as ’anxious’, ’happy’, ’optimistic’ and sentiment polarity prediction, such as positive, negative, or neutral opinions. Sentiment analysis has extensive applications in different domains, encompassing social media analytics, public sentiment monitoring, product review evaluation, and customer feedback assessment (Dang et al.,, 2020; Medhat et al.,, 2014). Recurrent neural networks (RNNs) (Salehinejad et al.,, 2017) are deep learning models that possess the ability to acquire features and capture long-term dependencies from sequential and time-series data (Salehinejad et al.,, 2017). The Transformer model is a deep learning model that features an encoder-decoder architecture and attention-based mechanism inspired by cognitive science (Vaswani et al.,, 2017). Although RNNs have been widely used, the introduction of the Transformer model has marked a significant shift in NLP, paving the way for pre-trained language models. Recent innovations in pre-trained language models, also known as language models (Kasneci et al.,, 2023), include BERT (Bidirectional Encoder Representation from Transformers) (Devlin et al.,, 2018) and GPT (Generative Pre-trained Transformers) (Kasneci et al.,, 2023), which have become prominent in a wide range of NLP tasks, exemplified by implementations like GPT-4o, Gemini 2.5, and Grok 3.
@@
Abuse detection using language models has been prominent due to the needs of social media, where real-time management of discussions are needed to avoid cyber-bullying. Caselli et al. (Caselli et al.,, 2020) introduced the Hate-BERT model for abusive content detection, achieving better accuracy in detecting various abusive content categories including offensive content, abusive content, and hateful language. Nobata et al. (Nobata et al.,, 2016) effectively developed a machine learning-based approach to detect hate speech in online user comments. Mozafari et al. (Mozafari et al.,, 2020) added a Bidirectional-LSTM layer to BERT’s architecture to enhance its classification capabilities. Mnassri et al. (Mnassri et al.,, 2022) explored various integrated approaches to improve BERT for hate speech detection, and reported that combining BERT with other deep learning models improved accuracy. Although these studies demonstrate the effectiveness of language models for abuse detection in social media contexts, their application to long-term cinematic data remains underexplored.
@@
The defining characteristic of longitudinal analysis lies in the repeated measurements taken on the same individuals or subjects of study over time, enabling a direct examination of temporal changes (Fitzmaurice et al.,, 2012). The primary objective of a longitudinal study is to delineate the dynamics of response variation and identify influential factors (Fitzmaurice et al.,, 2012). One can effectively examine temporal variations (Fitzmaurice et al.,, 2012) by capturing within-individual changes through multiple assessments. This motivates our study, which combines abuse detection and sentiment analysis to examine temporal trends in cinematic dialogues over more than seventy years.
@@
In this study, we present a framework for abuse detection and sentiment analysis on dialogues from selected movies. Hollywood movies (films) spanning from 1950 to 2024 were selected based on Oscar nominations and box-office collections. We analyse decadal changes in the levels of abuse and violence, utilising sentiment analysis across major categories, including Oscar nominations and blockbuster hits. We classify the selected movies into four categories to provide an analysis of how the abuse and sentiment analysis impact different categories of movies, e.g. action vs drama. An abuse detection dataset (Caselli et al.,, 2020) is used to fine-tune. BERT-based models and provide a longitudinal study. The framework also provides sentiment analysis, with the language models fine-tuned using the SenWave (Yang et al.,, 2020) dataset. The subtitle dataset of more than a thousand movies was also curated for further analysis.
This study addresses several gaps in the existing literature. First, prior studies on sentiment and abuse analysis have primarily focused on social media or reviews, with limited longitudinal analysis of cinematic dialogues across decades. Second, most studies focus on a single dimension (either sentiment or abusive content), this work integrates both to provide a more holistic perspective. Third, previous studies have not compared Oscar-nominated and blockbuster films and examined cross-genre differences. The contributions of this study are threefold:
(i) Curation and analysis of a large-scale subtitle corpus covering more than seven decades of Hollywood films.
(ii) Fine-tuning of advanced language models for simultaneous sentiment and abuse detection.
(iii) Provision of comparative insights across genres and award/commercial classifications, highlighting temporal shifts in cinematic language use.
@@
The structure of the paper is as follows: Section 2 provides background information, and Section 3 discusses the methodology, including the framework, data processing, and modelling. In Section 4, the results are presented. Section 5 offers a discussion and outlines directions for further research. Finally, Section 6 concludes the study.
@@ 2.1 Cinema: impact on society
Film as a medium is not only considered one of the most important sources of entertainment, but also a powerful tool for influencing society and culture (Saeed et al.,, 2018). Although the themes of love, friendship, and family relationships in films inspire audiences positively, some films with explicit, vulgar, or violent content can negatively impact viewers’ psychological well-being (Worth et al.,, 2008), affecting children and adolescents (Agarwal and Dhanasekaran,, 2012) . Furthermore, the characteristics of the protagonist, such as justice, wisdom and bravery, have a positive influence on the growth of young people (Smithikrai,, 2016).
The United Nations Educational, Scientific and Cultural Organisation (UNESCO) (Groebel,, 1998) presented a comprehensive study of the impact of media violence on the audience. Most studies show that the relationship between media abuse and “real” abuse is interactive, and the perpetuation and persistence of media violence fuelling a global culture of aggression (Ybarra et al.,, 2008; Morahan-Martin,, 1999; Mishna et al.,, 2009). An empirical study (Ferguson,, 2015) in 2015 challenged simplistic causal assumptions, showing that the relationship between media violence and societal aggression is highly context-dependent and varies across time periods. The American psychologist, Albert Bandura (Bandura and Walters,, 1977) reported that when people consciously or unconsciously remember and preserve abusive content, and unconsciously activate it, there are setbacks in social interactions. According to the American Academy of Pediatrics (on Communications and Media,, 2009), television violence accounts for 10 percent of the sources of learning about violent behaviour in children. Vitasari (Vitasari,, 2013) reported that the use of abusive words is now more often used in films. Using a corpus of film dialogues, Pavesi and Formentelli (Pavesi and Formentelli,, 2023) found that swearing serves coherence and emotional emphasis in audiovisual translation, rather than mere profanity. The uninhibited display of emotions is an indication of cultural transformation, resulting in a rise in the utilisation of profane language. Hasibuan (Hasibuan and Musfiroh,, 2021) contended that many films now often use swearing in dialogue to make them more interesting. Novari (Novari and Khoerunnisa,, 2022) reported that characters in movies do not always use swearing to curse but to express their other emotions and feelings in different situations. An editorial synthesis (Stapleton and Fägersten,, 2023) argued that swearing functions as a pragmatic strategy to convey interpersonal nuance beyond explicit meaning. Baid et al. (Baid et al.,, 2017) demonstrated that morphing learning from the social media domain can effectively classify hateful and offensive speech in movie subtitles. The authors used domain-adaptive and fine-tuned transfer learning techniques on an existing social media dataset. Deep learning models have been prominent in sentiment analysis (Chandra et al.,, 2023); hence, the study use them to analyse movie dialogues.
@@ 2.2 Web series and movies
Web series (Christian,, 2018) is a short-form interstitial content distributed through online platforms. Since the 1990s, with the advent of technological change and the emergence of the new media market, web series have become prominent, splitting the market from limited television and film (Christian,, 2012). During COVID-19, as countries imposed lockdowns and economic activity slowed, online sketches gained a significant opportunity to grow and reach more people. Traditional movies have been facing a huge challenge due to OTT platforms, and a study shows that as the length of movies increases, people expect movies released in theatres to be released simultaneously on OTT as well (Yan et al.,, 2021). Therefore, web series have become one of the most popular trends on OTT platforms (Deb,, 2022). However, the content preferences for films and web series are very different. In a related study, Wadhwa et al. (Wadhwa et al.,, 2020) conducted sentiment analysis for ten films of different genres as well as ten web series during 2017-2019. The authors accessed Twitter data related to each show through hashtags, and 3000 tweets were counted to determine the sentiment of each web series and categorised into positive and negative sentiment. The results found that horror web series were more popular with viewers than horror films, and categories such as comedy, romance and crime films were more popular than web series.
@@ 3.1 Data
The Oscars (Academy Awards)  [^1] Oscar website: https://www.oscars.org/is presented annually by the Academy of Motion Picture Arts and Sciences (AMPAS) and primarily recognises artistic and technical achievements in Hollywood movies (O’KEEFFE,, 2018). Historically, the Oscars have nominated a wide range of films, including historical dramas, romances, and science fiction. Although the drama category has traditionally dominated the Academy Awards, horror, comedy, and independent films have gradually gained more recognition in recent years with movies such as ”Get Out” and ”Backstage.” This diversity allows Oscar-nominated films to represent different periods and audiences’ aesthetic tastes (O’KEEFFE,, 2018). The annual top 10 global box office films represent the most popular movie genres and themes. These films typically attract large audiences and exert wide global influence. They span genres such as action, science fiction, comedy, and animation. Their global reception offers insights into cultural preferences and emotional resonance. The popularity of these films in different cultural contexts reflects the general acceptance of their emotional expressions and scenes of violence. Therefore, we focused on the subtitles of Oscar-nominated movies and the top 10 movie subtitles at the box office each year from 1950 to 2024.
@@
We obtained the top 10 Oscar-nominated movies and top 10 movies at the global Box Office for 1950 to 2024 from a website known as boxofficemojo  [^2] https://www.boxofficemojo.com/. We extracted the subtitles (dialogues) for the top 10 Oscar-nominated and global Box Office films from 1950 to 2024 from Opensubtitles  [^3] https://www.opensubtitles.com/en. The number of Oscar-nominated films varies across years. In total, we collected 1026 film subtitles, grouped by year and genre to analyse temporal trends in violence, abusive content, and emotional expression in films. The dataset is available via Kaggle  [^4] https://www.kaggle.com/datasets/mlopssss/subtitles and also documented in our GitHub repository  [^5] https://github.com/sydney-machine-learning/sentimentanalysis-Hollywood for this study.
@@
We classify the movies into four broad groups based on IMDb (International Movie Database)  [^6] https://www.imdb.com/ definitions, merging similar genres for comparison. We merged action, adventure, and crime into a single category called ”Action”. We combineed Horror and thriller into the ”Thriller” category. We keep the ”Comedy” and ”Drama” categories unchanged. Since most movies are defined by three category labels on the IMBb website, we take the first label as the main category of the movie as the basis for classification and reports the number of movies under each category in Table 1.
In addition to movie subtitles, we incorporated external datasets to train sentiment and abuse detection models. The SenWave (Yang et al.,, 2020) dataset for sentiment analysis comprises over 105 million collected tweets worldwide, providing an assessment of global sentiment fluctuations during the COVID-19 pandemic. This comprehensive dataset includes annotations for 10,000 English and 10,000 Arabic tweets across ten distinct categories: optimism, gratitude, compassion, pessimism, anxiety, sadness, anger, denial, official reports, and jokes. Qiang et al. (Loper and Bird,, 2002) preprocessed the raw data by filtering irrelevant characters using the NLTK (Natural Language Toolkit) tool. The SenWave dataset has been used to identify post-COVID-19 sentiment trends (Chandra and Krishna,, 2021) and to refine pre-trained language models for machine translation in low-resource languages (Chandra and Kulkarni,, 2022). These studies motivate the usage of the SenWave data for the analysis of Hollywood movie dialogues.
RAL-E (Reddit Abusive Language English) (Caselli et al.,, 2020) is a large-scale dataset comprising 1,478,348 messages from the Reddit 777www.blackdit.com online communities called Subredits (Ballinger,, 2022) that were banned for offensive, hateful, and abusive content. Subreddits were created and controlled by users, with moderators setting community rules (Matias,, 2016). In June 2015, Reddit banned hate speech-related Subreddits for violating the site’s user agreement and updated its content policy  [^8] https://www.redditinc.com/policies/content-policy. Tommaso et al. (Caselli et al.,, 2020) analysed the abusive vocabulary and compiled a data set for further research. The RAL-E data set has been used to train several transformer-based deep learning models, which have shown impressive performance in detecting online abuse (Caselli et al.,, 2020; Huang et al.,, 2022; Dacon et al.,, 2022). This study uses the RAL-E dataset to fine-tune the HateBERT model and improve its abuse detection capabilities.
@@
Table 1: Number of movies in the four categories that include drama, action, comedy and thriller.